# Unifying Unsupervised Graph-Level Anomaly Detection and Out-of-Distribution Detection: A Benchmark

### v1

@@ Abstract
To build safe and reliable graph machine learning systems, unsupervised graph-level anomaly detection (GLAD) and unsupervised graph-level out-of-distribution (OOD) detection (GLOD) have received significant attention in recent years. Though those two lines of research indeed share the same objective, they have been studied independently in the community due to distinct evaluation setups, creating a gap that hinders the application and evaluation of methods from one to the other. To bridge the gap, in this work, we present a Unified Benchmark for unsupervised Graph-level OOD and anomaLy Detection (UB-GOLD), a comprehensive evaluation framework that unifies GLAD and GLOD under the concept of generalized graph-level OOD detection. Our benchmark encompasses 35 datasets spanning four practical anomaly and OOD detection scenarios, facilitating the comparison of 16 representative GLAD/GLOD methods. We conduct multi-dimensional analyses to explore the effectiveness, generalizability, robustness, and efficiency of existing methods, shedding light on their strengths and limitations. Furthermore, we provide an open-source codebase (https://github.com/UB-GOLD/UB-GOLD) of UB-GOLD to foster reproducible research and outline potential directions for future investigations based on our insights.
@@ 1 Introduction
With the ubiquity of graph data, graph machine learning has been widely adopted in various scientific and industrial fields, ranging from bioinformatics to social networks [1; 2]. As one of the representative graph learning tasks, graph-level anomaly detection (GLAD) has been widely studied to identify abnormal graphs that show significant dissimilarity from the majority of graphs in a collection [3; 4]. GLAD task is crucial in various real-world applications, such as toxic molecule recognition and pathogenic brain mechanism discovery [5; 6]. Due to the high cost of data labeling, existing GLAD studies generally follow an unsupervised paradigm, eliminating the requirement of labeled anomaly samples for model training [3; 7; 8].
@@
In the meantime, another line of research – graph-level out-of-distribution detection (GLOD) – has drawn increasing attention in the research community lately [9; 10]. GLOD aims to identify whether each graph sample in a test set is in-distribution (ID), meaning it comes from the same distribution as the training data, or out-of-distribution (OOD), indicating it comes from different distributions. Considering the universality of distribution shift in open-world data, GLOD plays an important role in real-world high-stakes applications such as drug discovery and cyber-attack detection [11; 12; 13]. Though a few post-hoc GLOD methods [14; 15] that rely on a well-trained graph classifier have been proposed, most of the existing GLOD methods [9; 10; 11] are developed in an unsupervised fashion, where a specific OOD detection model is trained on unlabeled ID data, and then predicts a score for each test sample to indicate its ID/OOD status. In essence, unsupervised GLOD and unsupervised GLAD share the same goal, suggesting that research in one area could potentially be applied to solve problems in the other. Recent survey and benchmark papers [16; 17] also identify anomaly detection and OOD detection as two branches under the concept of “generalized OOD detection”, which further highlights the close conceptual connection between GLOD and GLAD tasks.
@@
Despite the inherent connection between the problems of unsupervised GLAD and GLOD, those two research sub-areas have been studied independently in the literature and have distinct evaluation setups. To fill the gap, in this paper, we develop a Unified Benchmark for unsupervised Graph-level OOD and anomaLy Detection (UB-GOLD for short). As shown in Fig. 1, we unify GLAD and GLOD as “generalized graph-level OOD detection problem”, and consider two GLOD scenarios and two GLAD scenarios for comprehensive evaluation. For GLAD task that aims to detect graph samples with semantic abnormality, we explore scenarios with intrinsic anomaly and class-based anomaly. Specifically, the first type of dataset inherently includes semantic anomalies, while the second type treats samples from one class (usually the minority class) as anomalies. For GLOD task that emphasizes the distribution shift between ID and OOD samples, we consider two scenarios with different distribution shifts, i.e., inter-dataset shift and intra-dataset shift. Specifically, the former simulates distribution shifts by drawing ID and OOD samples from different datasets within the same domain [9; 18], while the latter refers to datasets with intrinsic distribution shifts regarding graph sizes or molecular scaffolds [19]. Based on 35 datasets belonging to 4 types, we establish a comprehensive benchmark to fairly compare 16 GLAD and GLOD methods under a unified experimental setting, exploring the effectiveness of state-of-the-art (SOTA) approaches across diverse scenarios and domains.
@@
Apart from comparing performance, we further investigate the characteristics of GLAD/GLOD methods in terms of three dimensions: generalizability, robustness, and efficiency, through extensive experiments. For generalizability, we test a well-trained GLAD/GLOD model on OOD data with varying degrees of distribution shift, investigating each method’s ability to identify both near-OOD and far-OOD samples. For robustness, we introduce varying proportions of OOD samples into the training data to observe the impact of noisy ID data on the performance of existing methods. For efficiency, we perform efficiency evaluations for representative GLAD/GLOD approaches, focusing on time and space complexity. At the end of this paper, we discuss the future directions of this emerging research direction. To sum up, the main contribution of this paper is three-fold:
@@
Comprehensive benchmark. We introduce UB-GOLD, the first comprehensive and unified benchmark for GLAD and GLOD. UB-GOLD compares 16 representative GLAD/GLOD methods across 35 datasets from four practical anomaly and OOD detection scenarios.
@@
Multi-dimensional analysis. To explore their ability and limitations, we conduct a systematic analysis of existing methods from multiple dimensions, encompassing effectiveness on different datasets, generalizability to near-OOD and far-OOD, robustness against unreliable training data, and efficiency in terms of time and memory usage.
@@
Key findings. Through comprehensive comparison and analyses, we have the following remarkable observations: 1) The SOTA GLAD/GLOD methods show excellent performance across tasks; 2) Near-OOD samples are harder to detect compared to far-OOD samples; 3) Most methods are vulnerable to the perturbations of training sets; and 4) Certain end-to-end methods outperform 2-step methods in terms of both performance and computational costs.
@@ 2 Related Work and Problem Definition
Graph-level anomaly detection (GLAD). GLAD aims to identify abnormal graphs that show significant dissimilarity from the majority of graphs in a collection. A simple solution for GLAD is to use graph kernels [20; 21] to extract graph-level features and utilize shallow anomaly detectors (e.g. OCSVM [22] and iForest [23]) to identify anomalies. Recently, GNN-based end-to-end methods have demonstrated significant performance improvements in GLAD [3; 24]. Among them, some approaches assume that the anomaly labels of graphs are available for model training, and hence formulate GLAD as a supervised classification problem [4; 25; 26]. Nevertheless, their heavy reliance on labeled anomalies makes it challenging to apply them to real-world scenarios where annotated anomalies are scarce or unavailable. To overcome this shortage, the majority of GNN-based GLAD methods focus on an unsupervised learning paradigm, learning the GLAD model only from normal data [3; 7; 8; 24; 27; 28]. Recent research [9; 10] indicates that GLAD methods also show great potential in handling GLOD problem, prompting us to unify these two fields into a generalized research problem and conduct a comprehensive benchmarking study.
@@
Graph-level out-of-distribution detection (GLOD). GLOD refers to the problem of identifying whether a graph sample in a test set is ID or OOD, i.e., whether it originates from the same distribution as the training data or from a different one [10]. A line of studies developed post-hoc OOD detectors that identify OOD samples by additional fine-tuned detectors on top of well-trained GNN classifiers [14; 15]. However, these methods require annotated ID data to train the backbone GNNs, which limits their applicability in scenarios where labeled data is unavailable. In contrast, another line of research proposes training an OOD-specific GNN model using only ID data, without relying on any labels or OOD data [11; 9; 10]. To learn discriminative patterns of ID data, they employ unsupervised learning techniques such as contrastive learning [9] and diffusion model [11].
@@
Problem definition. Inspired by generalized OOD detection framework in computer vision [17; 19], we unify GLAD and GLOD into a high-level topic termed generalized graph-level OOD detection. The new learning task aims to distinguish generalized OOD samples (i.e., OOD or abnormal graphs) from generalized ID samples (i.e., ID or normal graphs). In this paper, we focus on the unsupervised scenario, considering its universality. The research problem is formulated as:
Definition 1 (Unsupervised generalized graph-level OOD detection).
@@
We define the training dataset as , where each sample is a graph drawn from a specific distribution . We define the testing dataset as , where contains graphs sampled from , and consists of graphs drawn from an OOD distribution . Given and , the learning objective is to train a detection model on and then the model can predict whether each sample belongs to or . In practice, is a scoring function, where a larger OOD score indicates a higher probability that is from (i.e., abnormal samples).
@@ 3 Benchmark Design
Dataset Type | Full Name | Abbreviation | Domain | OOD Definition | # ID Train | # ID Test | # OOD Test |
Tox21_p53 | p53 | Molecules | Inherent Anomaly | 8088 | 241 | 28 | |
(Type I) | Tox21_HSE | HSE | Molecules | Inherent Anomaly | 423 | 257 | 10 |
Intrinsic Anomaly | Tox21_MMP | MMP | Molecules | Inherent Anomaly | 6170 | 200 | 38 |
Tox21_PPAR-gamma | PPAR | Molecules | Inherent Anomaly | 219 | 252 | 15 | |
COLLAB | - | Social Networks | Unseen Classes | 1920 | 480 | 520 | |
IMDB-BINARY | IMDB-B | Social Networks | Unseen Classes | 400 | 100 | 100 | |
REDDIT-BINARY | REDDIT-B | Social Networks | Unseen Classes | 800 | 200 | 200 | |
ENZYMES | - | Proteins | Unseen Classes | 400 | 100 | 20 | |
(Type II) | PROTEINS | - | Proteins | Unseen Classes | 360 | 90 | 133 |
Class-based Anomaly | DD | - | Proteins | Unseen Classes | 390 | 97 | 139 |
BZR | - | Molecules | Unseen Classes | 69 | 17 | 64 | |
AIDS | - | Molecules | Unseen Classes | 1280 | 320 | 80 | |
COX2 | - | Molecules | Unseen Classes | 81 | 21 | 73 | |
NCI1 | - | Molecules | Unseen Classes | 1646 | 411 | 411 | |
DHFR | - | Molecules | Unseen Classes | 368 | 93 | 59 | |
IMDB-MULTI→IMDB-BINARY | IM→IB | Social Networks | Unseen Datasets | 1350 | 150 | 150 | |
ENZYMES→PROTEINS | EN→PR | Proteins | Unseen Datasets | 540 | 60 | 60 | |
AIDS→DHFR | AI→DH | Molecules | Unseen Datasets | 1800 | 200 | 200 | |
BZR→COX2 | BZ→CO | Molecules | Unseen Datasets | 364 | 41 | 41 | |
(Type III) | ESOL→MUV | ES→MU | Molecules | Unseen Datasets | 1015 | 113 | 113 |
Inter-Dataset Shift | TOX21→SIDER | TO→SI | Molecules | Unseen Datasets | 7047 | 784 | 784 |
BBBP→BACE | BB→BA | Molecules | Unseen Datasets | 1835 | 204 | 204 | |
PTC_MR→MUTAG | PT→MU | Molecules | Unseen Datasets | 309 | 35 | 35 | |
FREESOLV→TOXCAST | FS→TC | Molecules | Unseen Datasets | 577 | 65 | 65 | |
CLINTOX→LIPO | CL→LI | Molecules | Unseen Datasets | 1329 | 148 | 148 | |
GOOD-HIV-Size | HIV-Size | Molecules | Size | 1000 | 500 | 500 | |
GOOD-ZINC-Size | ZINC-Size | Molecules | Size | 1000 | 500 | 500 | |
GOOD-HIV-Scaffold | HIV-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | |
GOOD-ZINC-Scaffold | ZINC-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | |
(Type IV) | DrugOOD-IC50-Size | IC50-Size | Molecules | Size | 1000 | 500 | 500 |
Intra-Dataset Shift | DrugOOD-EC50-Size | EC50-Size | Molecules | Size | 1000 | 500 | 500 |
DrugOOD-IC50-Scaffold | IC50-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | |
DrugOOD-EC50-Scaffold | EC50-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | |
DrugOOD-IC50-Assay | IC50-Assay | Molecules | Protein Target | 1000 | 500 | 500 | |
DrugOOD-EC50-Assay | EC50-Assay | Molecules | Protein Target | 1000 | 500 | 500 |

### v2

@@ Abstract
To build safe and reliable graph machine learning systems, unsupervised graph-level anomaly detection (GLAD) and unsupervised graph-level out-of-distribution (OOD) detection (GLOD) have received significant attention in recent years. Though those two lines of research indeed share the same objective, they have been studied independently in the community due to distinct evaluation setups, creating a gap that hinders the application and evaluation of methods from one to the other. To bridge the gap, in this work, we present a Unified Benchmark for unsupervised Graph-level OOD and anomaLy Detection (UB-GOLD), a comprehensive evaluation framework that unifies GLAD and GLOD under the concept of generalized graph-level OOD detection. Our benchmark encompasses 35 datasets spanning four practical anomaly and OOD detection scenarios, facilitating the comparison of 18 representative GLAD/GLOD methods. We conduct multi-dimensional analyses to explore the effectiveness, OOD sensitivity spectrum, robustness, and efficiency of existing methods, shedding light on their strengths and limitations. Furthermore, we provide an open-source codebase (https://github.com/UB-GOLD/UB-GOLD) of UB-GOLD to foster reproducible research and outline potential directions for future investigations based on our insights.
@@ 1 Introduction
With the ubiquity of graph data, graph machine learning has been widely adopted in various scientific and industrial fields, ranging from bioinformatics to social networks (Xia et al., 2021; Wu et al., 2020; Zhang et al., 2024b; Wang et al., 2024b; Zhang et al., 2024a; Wang et al., 2022; Juan et al., 2024; Miao et al., 2024; Liu et al., 2024; Ding et al., 2024a). As one of the representative graph learning tasks, graph-level anomaly detection (GLAD) has been widely studied to identify abnormal graphs that show significant dissimilarity from the majority of graphs in a collection (Ma et al., 2022; Zhang et al., 2022). GLAD task is crucial in various real-world applications, such as toxic molecule recognition and pathogenic brain mechanism discovery (Jiang et al., 2021; Liu et al., ). Due to the high cost of data labeling, existing GLAD studies generally follow an unsupervised paradigm, eliminating the requirement of labeled anomaly samples for model training (Ma et al., 2022; Qiu et al., 2022; Liu et al., 2023b; Pan et al., 2023).
@@
In the meantime, another line of research – graph-level out-of-distribution detection (GLOD) – has drawn increasing attention in the research community lately (Liu et al., 2023a; Li et al., 2022; Wu et al., 2024). GLOD aims to identify whether each graph sample in a test set is in-distribution (ID), meaning it comes from the same distribution as the training data, or out-of-distribution (OOD), indicating it comes from different distributions. Considering the universality of distribution shift in open-world data, GLOD plays an important role in real-world high-stakes applications such as drug discovery and cyber-attack detection (Shen et al., 2024; Ji et al., 2023; Ju et al., 2024; Ding et al., 2024b; Li et al., 2024). Though a few post-hoc GLOD methods (Guo et al., 2023; Wang et al., 2024a) that rely on a well-trained graph classifier have been proposed, most of the existing GLOD methods (Liu et al., 2023a; Li et al., 2022; Shen et al., 2024) are developed in an unsupervised fashion, where a specific OOD detection model is trained on unlabeled ID data, and then predicts a score for each test sample to indicate its ID/OOD status. In essence, unsupervised GLOD and unsupervised GLAD share the same goal, suggesting that research in one area could potentially be applied to solve problems in the other. Recent survey and benchmark papers (Yang et al., 2022; 2021) also identify anomaly detection and OOD detection as two branches under the concept of “generalized OOD detection”, which further highlights the close conceptual connection between GLOD and GLAD tasks.
@@
Despite the inherent connection between the problems of unsupervised GLAD and GLOD, those two research sub-areas have been studied independently in the literature and have distinct evaluation setups. To fill the gap, in this paper, we develop a Unified Benchmark for unsupervised Graph-level OOD and anomaLy Detection (UB-GOLD for short). As shown in Fig. 1, we unify GLAD and GLOD as “generalized graph-level OOD detection problem”, and consider two GLOD scenarios and two GLAD scenarios for comprehensive evaluation. For GLAD task that aims to detect graph samples with semantic abnormality, we explore scenarios with intrinsic anomaly and class-based anomaly. Specifically, the first type of dataset inherently includes semantic anomalies, while the second type treats samples from one class (usually the minority class) as anomalies. For GLOD task that emphasizes the distribution shift between ID and OOD samples, we consider two scenarios with different distribution shifts, i.e., inter-dataset shift and intra-dataset shift. Specifically, the former simulates distribution shifts by drawing ID and OOD samples from different datasets within the same domain (Liu et al., 2023a; Sehwag et al., 2021), while the latter refers to datasets with intrinsic distribution shifts regarding graph sizes or molecular scaffolds (Gui et al., 2022). Based on 35 datasets belonging to 4 types, we establish a comprehensive benchmark to fairly compare 18 GLAD and GLOD methods under a unified experimental setting, exploring the effectiveness of state-of-the-art (SOTA) approaches across diverse scenarios and domains.
@@
Apart from comparing performance, we further investigate the characteristics of GLAD/GLOD methods in terms of three dimensions: OOD sensitivity spectrum, robustness, and efficiency, through extensive experiments. For OOD sensitivity spectrum, we test a GLAD/GLOD model on OOD data with varying degrees of distribution shift, investigating each method’s ability to identify both near-OOD and far-OOD samples. For robustness, we introduce varying proportions of OOD samples into the training data to observe the impact of noisy ID data on the performance of existing methods. For efficiency, we perform efficiency evaluations for representative GLAD/GLOD approaches, focusing on time and space complexity. At the end of this paper, we discuss the future directions of this emerging research direction. To sum up, the main contribution of this paper is three-fold:
@@
Comprehensive benchmark. We introduce UB-GOLD, the first comprehensive and unified benchmark for GLAD and GLOD. UB-GOLD compares 18 representative GLAD/GLOD methods across 35 datasets from four practical anomaly and OOD detection scenarios.
@@
Multi-dimensional analysis. To explore their ability and limitations, we conduct a systematic analysis of existing methods from multiple dimensions, encompassing effectiveness on different datasets, OOD sensitivity spectrum to near-OOD and far-OOD, robustness against unreliable training data, and efficiency in terms of time and memory usage.
@@
Key findings. Through comprehensive comparison and analyses, we have the following remarkable observations: 1) The SOTA GLAD/GLOD methods show excellent performance across tasks; 2) Near-OOD samples are harder to detect compared to far-OOD samples; 3) Most methods are vulnerable to the perturbations of training sets; and 4) Certain end-to-end methods outperform two-step methods in terms of both performance and computational costs.
@@ 2 Related Work and Problem Definition
Graph-level anomaly detection (GLAD). GLAD aims to identify abnormal graphs that show significant dissimilarity from the majority of graphs in a collection. A simple solution for GLAD is to use graph kernels (Vishwanathan et al., 2010; Neumann et al., 2016) to extract graph-level features and utilize shallow anomaly detectors (e.g. OCSVM (Amer et al., 2013) and iForest (Liu et al., 2008)) to identify anomalies. Recently, GNN-based end-to-end methods have demonstrated significant performance improvements in GLAD (Ma et al., 2022; Zhao & Akoglu, 2023). Among them, some approaches assume that the anomaly labels of graphs are available for model training, and hence formulate GLAD as a supervised classification problem (Zhang et al., 2022; Ma et al., 2023; Dong et al., 2024). Nevertheless, their heavy reliance on labeled anomalies makes it challenging to apply them to real-world scenarios where annotated anomalies are scarce or unavailable. To overcome this shortage, the majority of GNN-based GLAD methods focus on an unsupervised learning paradigm, learning the GLAD model only from normal data (Ma et al., 2022; Qiu et al., 2022; Liu et al., 2023b; Zhao & Akoglu, 2023; Luo et al., 2022; Li et al., 2023). Recent research (Liu et al., 2023a; Li et al., 2022) indicates that GLAD methods also show great potential in handling GLOD problem, prompting us to unify these two fields into a generalized research problem and conduct a comprehensive benchmarking study.
@@
Graph-level out-of-distribution detection (GLOD). GLOD refers to the problem of identifying whether a graph sample in a test set is ID or OOD, i.e., whether it originates from the same distribution as the training data or from a different one (Li et al., 2022). A line of studies developed post-hoc OOD detectors that identify OOD samples by additional fine-tuned detectors on top of well-trained GNN classifiers (Guo et al., 2023; Wang et al., 2024a). However, these methods require annotated ID data to train the backbone GNNs, which limits their applicability in scenarios where labeled data is unavailable. In contrast, another line of research proposes training an OOD-specific GNN model using only ID data, without relying on any labels or OOD data (Shen et al., 2024; Liu et al., 2023a; Li et al., 2022). To learn discriminative patterns of ID data, they employ unsupervised learning techniques such as contrastive learning (Liu et al., 2023a) and diffusion model (Shen et al., 2024).
@@
Problem definition. Inspired by generalized OOD detection framework in computer vision (Yang et al., 2021; Gui et al., 2022), and due to the highly similar objectives of GLAD and GLOD (for detailed discussion, see Appendix D.1), we unify GLAD and GLOD into a high-level topic termed generalized graph-level OOD detection. The new learning task aims to distinguish generalized OOD samples (i.e., OOD or abnormal graphs) from generalized ID samples (i.e., ID or normal graphs). In this paper, we focus on the unsupervised scenario, considering its universality. The research problem is formulated as:
@@
Definition 1. [Unsupervised generalized graph-level OOD detection] We define the training dataset as , where each sample is a graph drawn from a specific distribution . We define the testing dataset as , where contains graphs sampled from , and consists of graphs drawn from an OOD distribution . Given and , the learning objective is to train a detection model on and then the model can predict whether each sample belongs to or . In practice, is a scoring function, where a larger OOD score indicates a higher probability that is from (i.e., abnormal samples).
The theoretical objective of the scoring function is to maximize the separation between the distributions of ID/normal graphs and OOD/anomalous graphs:
This unified objective integrates the detection principles of GLAD and GLOD, emphasizing their shared goal of distinguishing graphs that deviate from a given distribution.
@@ 3 Benchmark Design
Dataset Type | Full Name | Abbreviation | Domain | OOD Definition | # ID Train | # ID Test | # OOD Test | # Anomaly Ratio % |
Tox21_p53 | p53 | Molecules | Inherent Anomaly | 8088 | 241 | 28 | 0.34 | |
(Type I) | Tox21_HSE | HSE | Molecules | Inherent Anomaly | 423 | 257 | 10 | 1.45 |
Intrinsic Anomaly | Tox21_MMP | MMP | Molecules | Inherent Anomaly | 6170 | 200 | 38 | 0.59 |
Tox21_PPAR-gamma | PPAR | Molecules | Inherent Anomaly | 219 | 252 | 15 | 3.09 | |
COLLAB | - | Social Networks | Unseen Classes | 1920 | 480 | 520 | 17.81 | |
IMDB-BINARY | IMDB-B | Social Networks | Unseen Classes | 400 | 100 | 100 | 16.67 | |
REDDIT-BINARY | REDDIT-B | Social Networks | Unseen Classes | 800 | 200 | 200 | 16.67 | |
ENZYMES | - | Proteins | Unseen Classes | 400 | 100 | 20 | 3.85 | |
(Type II) | PROTEINS | - | Proteins | Unseen Classes | 360 | 90 | 133 | 22.81 |
Class-based Anomaly | DD | - | Proteins | Unseen Classes | 390 | 97 | 139 | 22.20 |
BZR | - | Molecules | Unseen Classes | 69 | 17 | 64 | 42.67 | |
AIDS | - | Molecules | Unseen Classes | 1280 | 320 | 80 | 4.76 | |
COX2 | - | Molecules | Unseen Classes | 81 | 21 | 73 | 41.71 | |
NCI1 | - | Molecules | Unseen Classes | 1646 | 411 | 411 | 16.65 | |
DHFR | - | Molecules | Unseen Classes | 368 | 93 | 59 | 11.35 | |
IMDB-MULTI→IMDB-BINARY | IM→IB | Social Networks | Unseen Datasets | 1350 | 150 | 150 | 9.09 | |
ENZYMES→PROTEINS | EN→PR | Proteins | Unseen Datasets | 540 | 60 | 60 | 9.09 | |
AIDS→DHFR | AI→DH | Molecules | Unseen Datasets | 1800 | 200 | 200 | 9.09 | |
BZR→COX2 | BZ→CO | Molecules | Unseen Datasets | 364 | 41 | 41 | 9.19 | |
(Type III) | ESOL→MUV | ES→MU | Molecules | Unseen Datasets | 1015 | 113 | 113 | 9.11 |
Inter-Dataset Shift | TOX21→SIDER | TO→SI | Molecules | Unseen Datasets | 7047 | 784 | 784 | 9.10 |
BBBP→BACE | BB→BA | Molecules | Unseen Datasets | 1835 | 204 | 204 | 9.09 | |
PTC_MR→MUTAG | PT→MU | Molecules | Unseen Datasets | 309 | 35 | 35 | 9.23 | |
FREESOLV→TOXCAST | FS→TC | Molecules | Unseen Datasets | 577 | 65 | 65 | 9.19 | |
CLINTOX→LIPO | CL→LI | Molecules | Unseen Datasets | 1329 | 148 | 148 | 9.11 | |
GOOD-HIV-Size | HIV-Size | Molecules | Size | 1000 | 500 | 500 | 25.00 | |
GOOD-ZINC-Size | ZINC-Size | Molecules | Size | 1000 | 500 | 500 | 25.00 | |
GOOD-HIV-Scaffold | HIV-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | 25.00 | |
GOOD-ZINC-Scaffold | ZINC-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | 25.00 | |
(Type IV) | DrugOOD-IC50-Size | IC50-Size | Molecules | Size | 1000 | 500 | 500 | 25.00 |
Intra-Dataset Shift | DrugOOD-EC50-Size | EC50-Size | Molecules | Size | 1000 | 500 | 500 | 25.00 |
DrugOOD-IC50-Scaffold | IC50-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | 25.00 | |
DrugOOD-EC50-Scaffold | EC50-Scaffold | Molecules | Scaffold | 1000 | 500 | 500 | 25.00 | |
DrugOOD-IC50-Assay | IC50-Assay | Molecules | Protein Target | 1000 | 500 | 500 | 25.00 | |
DrugOOD-EC50-Assay | EC50-Assay | Molecules | Protein Target | 1000 | 500 | 500 | 25.00 |